### v1

# Hierarchical Scheduling for Multi-Vector Image Retrieval

Abstract
To effectively leverage user-specific data, retrieval-augmented generation (RAG) is employed in multimodal large language model (MLLM) applications. However, conventional retrieval approaches often suffer from limited retrieval accuracy. Recent advances in multi-vector retrieval (MVR) improve accuracy by decomposing queries and matching against segmented images. They still suffer from sub-optimal accuracy and efficiency, overlooking alignment between the query and varying image objects and redundant fine-grained image segments. In this work, we present an efficient scheduling framework for image retrieval – HiMIR. First, we introduce a novel hierarchical paradigm, employing multiple intermediate granularities for varying image objects to enhance alignment. Second, we minimize redundancy in retrieval by leveraging cross-hierarchy similarity consistency and hierarchy sparsity to minimize unnecessary matching computation. Furthermore, we configure parameters for each dataset automatically for practicality across diverse scenarios. Our empirical study shows that, HiMIR not only achieves substantial accuracy improvements but also reduces computation by up to $3.5\times$ over the existing MVR system.
footnotetext: $\dagger$ Equal Contribution. * Corresponding to Xiang Chen (xiang.chen@pku.edu.cn).
I Introduction
The rapid advancement of large language models (LLMs) has enabled emerging applications such as intelligent agents and personal assistants [1]. However, LLMs are not inherently capable of effectively leveraging user-specific data. To address this limitation, retrieval-augmented generation (RAG) has been widely adopted, where external data relevant to a user query is retrieved and incorporated into the generation process to improve accuracy and relevance [2]. Meanwhile, when RAG is extended to multimodal LLM systems, specifically with image retrieval, it computes the similarity between the feature vectors of language prompts and particular image objects [3]. Most multimodal RAG systems in production, however, follow a simple one-shot paradigm. As shown in Fig. 1, MVR embeds an entire query and an entire image into a single global vector, referred to as “1 Mode” in this context. While efficient, this single-vector retrieval inevitably loses fine-grained object information, leading to unsatisfactory retrieval accuracy for complex or semantically diverse image content.
To overcome this limitation, recent studies have explored retrieval with data decomposition. As illustrated in Fig. 1, another approach – multi-vector retrieval (MVR) decomposes a query into multiple independent sub-queries by semantic clustering with LLM prompts, while decomposing each image into N segments via granularity segmentation. Then, the sub-query embeddings match against N image segment embeddings. And the multi-vector retrieved information can be further aggregated with the single-vector retrieved information to balance global image and fine-grained object granularity. Such “1+N Mode” represents a fundamental shift toward better granularity of alignment between queries and the unpredictable semantic object composition of images.
Figure 1: Multimodal Image Retrieval
However, despite its accuracy benefits, MVR introduces several challenges in terms of efficiency and practicality. First, the number of decomposed vectors (N) is often set to be large in order to capture fine-grained objects. Yet, selecting the optimal N is highly non-trivial: too small an N fails to fully represent image granularity, while too large an N breaks the integrity of objects. Second, finer decomposition inherently amplifies computation complexity, as each additional image segment takes extra similarity calculation against multiple sub-queries. Finally, although finer-grained decomposition reveals structural properties such as redundancy across image segments and sparsity within query–image alignments, prior work has largely overlooked exploiting these opportunities [4, 5].
Figure 2: Image Retrieval with Hierarchical Decomposition
These challenges motivate the design of HiMIR, a retrieval scheduling framework that introduces hierarchical decomposition into multi-vector image retrieval:
From the algorithm perspective, unlike the conventional “1+N Mode” MVR, HiMIR extends it into a “1+M+N Mode” as shown in Fig. 2, where M represents a series of intermediate image segmentation granularities and hierarchical query matching. This hierarchy effectively adapts to images with varying object granularities, significantly improving alignment robustness and, therefore, retrieval accuracy.
From the computing perspective, although “1+M+N Mode” seems to further increase the computation complexity compared to “1+N” MVR, it actually exposes a more favorable computing optimization space. It exploits the cross-hierarchy consistency of retrieval information to reduce redundant matching, while simultaneously detecting and pruning sparsity per-hierarchy to minimize unnecessary computations.
From the design automation perspective, HiMIR further enhances practicality by supporting automatic configuration across datasets. Given the dataset variability, it conducts a lightweight profiling of dataset characteristics and derives the optimal parameterization for hierarchical decomposition. This enables joint automation of both algorithmic decomposition and computational scheduling, significantly improving adaptability across diverse deployment scenarios.
Following such a design methodology of HiMIR, this work made the following contributions:
  • We present a hierarchical decomposition framework for multi-vector image retrieval, offering a novel approach adaptive to diverse image granularities.
  • We systematically exploit sparsity in multi-vector retrieval, enabling a runtime acceleration mechanism that eliminates redundant computations.
  • We integrate these techniques into an automated framework that jointly optimizes accuracy and efficiency, enabling robust and adaptive deployment across scenarios.
[…]
II Background
II-A Decomposition in Retrieval
In multimodal RAG systems, the retrieval task aims to identify the most relevant image for a given input prompt query. Both the query and images are embedded into a shared vector space using text–image embedding models [6, 7], and the top-$K$ images with the highest similarity scores to the query are retrieved. The retrieval process can be expressed as Eq. 1, where $Q$ denotes the query, $D$ the image set, and $E(\cdot)$ the embedding model. $\operatorname{SIM}$ represents a similarity operator (e.g., cosine similarity, dot product, or L1- distance).
$\underset{0<i\leq N_{D}}{\operatorname{TopK}}(\operatorname{Score}(Q,D_{i}))=\operatorname{TopK}(\operatorname{SIM}(E(Q),E(D_{i})))$ (1)
However, encoding a semantically complex query or image into a single vector inevitably incurs information loss, resulting in sub-optimal retrieval quality. To address this issue, one-shot decomposition-based methods, or MVR (e.g., ColBERT and its variants [8, 9]), have been proposed. As formulated in Eq. 2, MVR decomposes a query into multiple sub-queries $q_{i}$} and each image into multiple segments $D_{i,j}$}. The overall score is computed as the product of the scores of all sub-queries, where the score of each sub-query is defined as the maximum similarity against all image segments. In other words, each sub-query matches its most relevant image segment, ensuring that all semantic components of the query are satisfied. To further improve retrieval accuracy, recent studies [4] adopt the “1+N Mode”, which aggregates scores computed with and without decomposition.
$\begin{split}\operatorname{Score}(Q,D_{i})=&\operatorname{SIM}(E(Q),E(D_{i}))\\ +&\prod_{i=1}^{N_{q}}\max_{1\leq j\leq m}\operatorname{SIM}(E(q_{i}),E(D_{i,j})).\end{split}$ (2)
Despite of its accuracy benefits, there is still ample space for improving MVR: First, granularity selection and alignment in image decomposition has not been systematically explored. Second, it introduces heavy overhead as similarity calculation increases by tens of times due to “N” matching.
II-B Decomposition Granularity
The decomposition granularity in MVR, i.e., the number of decomposed vectors N, plays a critical role in retrieval accuracy. Granularity can be considered from two perspectives:
Query decomposition granularity. A query should be decomposed into semantically independent sub-queries to achieve precise alignment with image objects. Coarse-grained decomposition often results in information loss, while overly fine-grained ones like token-level [8], may break semantic integrity and cause spurious matches. Recent work [4] leverages fine-tuned LLMs [10] to adaptively decompose complex queries.
Image decomposition granularity. Unlike text retrieval, where segmentation boundaries naturally exist at the sentence or paragraph level, image retrieval lacks inherent structural boundaries. Consequently, fast adaptive segmentation methods such as SLIC (Simple Linear Iterative Clustering) [11] are employed to partition images into segments, with granularity typically set empirically in the range of 4 to 64.
The major limitation of existing image decomposition approaches is that granularity is empirically pre-defined and lacks runtime adaptability.
II-C Approximate Retrieval Acceleration
As the complexity and scale of retrieval grow, computational redundancy becomes non-trivial, motivating the adoption of approximate methods that trade accuracy for efficiency. Most existing approximate acceleration techniques are algorithm-agnostic: indexing-based methods, such as IVF[12], HNSW [13], and PLAID [14], reduce the number of retrieval candidates for each query via clustering embedding vectors; quantization-based approaches such as IVF-PQ [12] reduce similarity computation latency by lowering data bit-width, thereby enhancing hardware parallelism. While, our approach exploits redundancy inherent in the MVR algorithm itself, which is largely orthogonal to these techniques and can thus be seamlessly integrated into existing optimizations.
III Hierarchical Decomposition
In this section, we present the analysis for the mis-alignment between query and image segments regarding retrieval granularities. And we propose a novel hierarchical decomposition algorithm framework, which both enables better alignment and exposes new scheduling opportunities.
Figure 3: Decomposition Granularity and Accuracy Impact
III-A Granularity Misalignment Analysis
Though the “1+N Mode” achieves non-trivial accuracy improvements, it still suffers from misalignment between fixed decomposition granularity and the variable scales of image objects. As shown in Fig. 3(a), which visualizes an image (same as in Fig. 5) decomposed into 25 segments with ground-truth objects marked by colored rectangles, such misalignment becomes evident. For certain objects, such as the corkboard (highlighted in red), decomposition is beneficial: irrelevant regions are cropped out while the main body of the object remains intact. However, other objects are adversely affected when the granularity is misaligned: objects may be fragmented into parts (e.g., the shelves marked in green) or merged with irrelevant regions (e.g., the toy marked in yellow).
To further prove this observation, we conduct a preliminary experiment by profiling accuracy across a set of granularities using the “1+N Mode” in Eq. 2. The results on two datasets are plotted with a solid line in Fig. 3(b). As the decomposition granularity becomes finer, the retrieval accuracy generally shows an upward trend and then levels off. However, accuracy decreases at both certain granularities in the middle and the finest ones. This can be attributed to the misalignment. Hence, to attain higher retrieval accuracy, we need to align the decomposition granularity “N” with each object. This could hardly be implemented with a single fixed granularity, so we turn to the aggregation of multiple granularities.
III-B Hierarchical Aggregation
Inspired by the analysis above, we want to align each required object with its best-match granularity for accuracy improvement. Thus, we extend the “1+N Mode” MVR algorithm by building the hierarchy of granularities, and searching along it for the most suitable granualrity of each sub-query (corresponding to an image object). The proposed algorithm can be denoted “1+M+N Mode” as illustrated in Fig. 2. Specifically, the images are segmented hierarchically into multiple granularities, each denoting a different patch size. The similarity score between each sub-query and image segments at a specific granularity is computed in the same way as Eq. 2. The final score for each sub-query is aggregated by selecting the maximum similarity score among all granularities, which represents the best alignment. With granularity denoted $g$, segment count in granularity $g$ denoted $N_{g}$, and the total number of granularities denoted $N_{G}$, we can extend Eq. 2 into Eq. 3:
$\operatorname{Score}(Q,D_{i})=\prod_{k=1}^{N_{q}}\max_{g=1}^{N_{G}}\max_{j=1}^{N_{g}}\operatorname{SIM}(E(q_{k}),E(D_{i,j}^{g}))$ (3)
Our “1+M+N Mode” is superior to the prevalent “1+N Mode” in two aspects. First, by considering all possible granularities, it adaptively leverages the most suitable segment size for each sub-query. As shown in Fig. 5, the sub-query “toy” matches the granularity of 9 while the other two objects match the granularity of 16. As shown in Fig. 3, the accuracy of our“1+M+N Mode” is always higher than “1+N Mode” and grows monotonically as the granularity becomes finer. Furthermore, our approach provides flexibility for an accuracy-performance trade-off by allowing for computing the relevance score on a selected subset of the hierarchy, which will be elaborated in the next section.
IV Computing Efficiency Exploration
The proposed hierarchical framework decouples the retrieval process into structured levels, enabling systematic analysis of information redundancy and principled optimization of computational efficiency.
Figure 4: Computing Redundancy Analysis
IV-A Redundancy Analysis with Retrieval Granularity
Figure 5: Computation Optimization in HiMIR
Deriving from the computation in Eq. 3, redundancy can be analyzed from two perspectives: the image dimension $D_{i}$ and the granularity dimension $D^{g}$.
Image dimension. Intuitively, coarse-grained image segments still preserve information about finer-grained objects, despite embedding loss. Neglecting the reuse of such information, which is consistent across the granularity hierarchy, leads to significant redundant computation. We validate this hypothesis by evaluating the top-$k$ recall and the rank distribution of ground-truth images across granularities, as shown in Fig. 4(a). The results indicate that ground-truth images consistently appear near the top ranks at all granularities. This observation suggests that even coarse-grained representations are sufficient to separate relevant images from irrelevant ones.
Granularity dimension. Although object sizes vary across images, they are not evenly distributed across all granularities. Exhaustively scanning every granularity for each query, therefore, introduces redundancy. To quantify this, we compute the best-matching granularity for each sub-query, and the distribution is illustrated in Fig. 4(b). The results reveal that certain granularities are rarely utilized, leading to unnecessary similarity computations for all queries. Moreover, the finest-grained representation is not always required, which further contributes to wasted computation in some cases.
Based on our analysis of granularity and redundancy, we identify three complementary dimensions of computational optimization within hierarchical decomposition. As illustrated in Fig. 5, these dimensions highlight distinct opportunities for reducing redundant computations and improving efficiency, which we elaborate on in the following subsections.
[…]
Image granularity analysis shows that the similarity score of ground-truth image consistently ranks near the top across all granularities. As more granularities are considered, the rank of the ground-truth image improves, whereas images ranked lower rarely benefit, even with finest-grained matching. This cross-hierarchy consistency suggests that performing fine-grained retrieval on low-ranking images is largely redundant. Based on this observation, we design an online optimization algorithm that tracks the score ranking of images in each granularity hierarchy. By eliminating costly fine-grained computations on low-confidence candidate images, the algorithm reduces runtime overhead without compromising accuracy.
Specifically, we compute similarity scores with image segments in each granularity hierarchy iteratively. In each iteration (hierarchy), images with low aggregated similarity scores are pruned and will not be computed in the following granularities. The procedure is elaborated in L.2–L.6 of Alg. 1. The ratio of images pruned in each granularity hierarchy is specified with $t_{g}=\alpha^{g}T$, where $T$ denotes the initial reduction ratio and $\alpha$ denotes the decay rate. Empirically, $T$ can be more aggressive, as ground truth usually ranks near the top. While it is better to set a conservative $\alpha$, in that the score rank of ground truth may fluctuate as the granularity becomes finer.
Algorithm 1 Online Redundancy Reduction
1:procedure ProcessQuery($\{q_{k}\}$)
2:Let $\operatorname{Score}[N_{D},N_{q},N_{G}]$, $\operatorname{TopK}[N_{G}]$
3:for $g:=1$ to $N_{G}$ do
4:  for $k\in[1,N_{q}]$, $i\in[1,N_{D}t_{g-1}]$, $j\in[1,N_{g}]$ do
5:   $\operatorname{Score}_{i,k,g}\leftarrow\max(\operatorname{SIM}(q_{k},D_{i,j}^{g}),\operatorname{Score}_{i,k,g-1})$
[…]
7:  $\operatorname{TopK}[g]:=D[1:N_{K}]$
8:  if $\operatorname{K\tau}(\text{TopK}[g],\text{TopK}[g-1])\geq\tau$ then
9:   break
10:return $\operatorname{TopK}[g]$
IV-C Early Retrieval Convergence $\rightarrow$ Hierarchy Depth Opt.
Hierarchy distribution analysis reveals that a substantial portion of queries do not align with the finest granularity. For these queries, matching against the finest-grained image segments is unnecessary and introduces excessive redundancy. To address this issue, we design a hierarchy skipping mechanism that monitors the confidence level across granularities. By terminating the search early once a confidence threshold is reached, the mechanism adaptively avoids redundant fine-grained computations, incurring only minimal accuracy loss.
The key challenge is to define the confidence indicator. According to Eq. 3, similarity scores remain unchanged until one of the sub-queries finds a better match. Therefore, it is natural to stop when the top-$k$ result list stabilizes. To quantify this convergence, we employ Kendall’s $\tau$ coefficient [15], denoted $\operatorname{K\tau}$, which measures the ordinal association between two ranked sequences. A preliminary study (Fig. 4(c)) confirms that $\operatorname{K\tau}$ consistently converges across datasets. In other words, when $\operatorname{K\tau}$ reaches a steady value, finer-grained segments no longer improve the result. The procedure is detailed in Alg. 1, Lines 7–-9, where $\textit{TopK}[g]$ denotes the set of $N_{K}$ images with the highest similarity scores up to granularity $g$.
IV-D Retrieval Granularity Sparisty $\rightarrow$ Hierarchy Sparisty Opt.
Beyond depth redundancy, cross-hierarchy analysis further indicates that certain intermediate granularities rarely align optimally with image objects. Thus they introduce additional redundancy, since it is impractical to make skipping choices online without prior knowledge of the dataset. Motivated by this insight, we propose an offline hierarchy pruning algorithm that eliminates redundant granularities based on dataset profiling. This reduces retrieval cost without sacrificing accuracy, since best-aligned objects of pruned granularities can typically be captured in adjacent ones.
Alg. 2 (L.1–L.12) details the algorithm. We construct the hierarchy by evenly distributing levels over $[1,N]$ with interval $S_{G}$, where a smaller $S_{G}$ yields higher flexibility but more redundancy. The algorithm iteratively removes the hierarchy with the least accuracy loss. To ensure scheduling flexibility, consecutive removals are prohibited to prevent scenarios where only fine-grained hierarchies remain, which limits opportunities for online optimizations.
V Framework Integration & Automation
Based on the redundancy theory introduced in Section IV, we implement the HiMIR framework with an automated configuration algorithm for optimizing parameters towards various trade-off objectives.
V-A Automated Configuration
HiMIR can flexibly adapt to different dataset characteristics and deployment scenarios, whether accuracy- or performance-oriented, by exposing a set of tunable parameters: the initial ratio $T$ and decay rate $\alpha$ for redundant matching optimization, the hierarchy skipping threshold $\tau$, and the granularity initialization stride $S_{G}$, as detailed in Section IV.
Efficiently optimizing over this high-dimensional parameter space is non-trivial. Thus, we design a latency-guided configuration algorithm based on grid search, as presented in Alg. 2. For each parameter, the search range and step size are predefined. During optimization, we traverse the latency dimension to limit profiling overhead, guided by a lightweight performance model:
$\operatorname{Latency}=N_{q}\sum\nolimits_{g=1}^{N_{G}}N_{g}\times t_{g}\times N_{D},$ (4)
where $N_{D}$ denotes the image set size, $N_{G}$, $N_{q}$, $N_{g}$ are the granularities, sub-queries, and the segments per granularity $g$, and $t_{g}$ is the fraction of images preserved at hierarchy $g$.
Since similarity computation is lightweight, the latency of matching each sub-query with an image segment can be treated as a constant and is omitted here for simplicity. Constrained by latency, we first establish the granularity hierarchy, which requires more extensive exploration (L.15). Subsequently, other parameters are tuned for the highest accuracy under each latency constraint as (L.16–L.17). The latency-guided search enables efficient automated configuration. With proper initialization, the entire procedure completes within tens of minutes on datasets in our experiments.
Algorithm 2 Automated Configuration
1:procedure SetGran($S_{G},\tau,\alpha,T$)
2:repeat
3:  $N_{g}:=N_{g}+S_{G};\ \{N_{g}\}:=\{N_{g}\}\cup N_{g}$
4:until $\text{Eval}(\{N_{g}\},\tau,\alpha,T)$ Converge to $\mathcal{A}$
5:repeat
6:  for $g:=1$ to $|\{N_{g}\}|$ do
7:   $\{Acc\}:=\{Acc\}\cup\text{Eval}(\{N_{g}\}-N_{g},\tau,\alpha,T)$
8:  repeat
9:   $g:=\operatorname{argmin}(\{\text{Acc}\})$; $\{\text{Acc}\}:=\{Acc\}-g$
10:  until $g$ not near the last removed granularity
11:  $\{N_{g}\}:=\{N_{g}\}-g$
12:until $\text{Eval}(\{N_{g}\})\leq\mathcal{A}$
13:procedure Configure($R_{\tau},R_{S},R_{\alpha},R_{T},R_{L}$)
14:for $L,S,\tau,\alpha,T$ in $R_{L},R_{S},R_{\tau},R_{\alpha},R_{T}$ do
15:  Let $\{N_{g}\}:=\textsc{SetGran}(S,R_{\tau},R_{\alpha},R_{T})$
16:  if $L\geq\operatorname{Latency}(\tau,T,\alpha,\{N_{g}\})$ and $\text{Cfg[L].Acc}<\operatorname{Eval}(\tau,T,\alpha,\{N_{g}\})$ then
17:   $\text{Cfg}[L]:=\{\tau,T,\alpha,\{N_{g}\}\}$
V-B Implementation Detail
In this section, we reveal details of our implementation. We added about 1k lines of code based on the open-source code from  [4] built on PyTorch [16] and Faiss [12]. For image decomposition, given that the boundaries of segments generated by SLIC are irregular, we pad the remaining area with a black background to the bounding box of each segment to form valid patches. For query decomposition, we deployed a local large language model using vLLM [17], where the prompt template was adopted directly from [4]. For query processing, we optimize performance by vectorizing and parallelizing CPU-GPU operations with minimal memory transferring overhead.
VI Experiments
TABLE I: Main Results
Datasets Crepe MS Coco NoCaps Flickr Average
NDCG@10 QPS NDCG@10 QPS NDCG@10 QPS NDCG@10 QPS Spd. Acc.
Vanilla 65.11 1230 66.27 678.1 82.10 638.4 82.30 601.4
POQD 71.62 54.8 68.44 27.8 84.68 27.8 85.19 29.7 1x 1x(+3.53)
Ours (w/o Opt.) 73.22 12.1 70.13 6.7 85.60 7.1 87.70 6.9 0.25x 1.48x(+5.22)
Ours (w/o O1,O2) 73.17 21.1 70.39 10.8 85.45 10.9 87.59 11.3 0.4x 1.47x(+5.2)
Ours (w/o O1) 73.56 133.3 70.60 78.1 85.84 76.23 87.24 72.8 2.5x 1.52x(+5.37)
Ours 72.83 148.6 70.27 93.3 85.23 101.3 86.57 102.9 3.5x 1.43x(+5.03)
VI-A Experiment Setup
Evaluation Platform. As our framework is hardware-agnostic, we evaluate it on a standard platform equipped with an Intel Xeon 4410T CPU and one NVIDIA A100-PCIE-40G GPU.
Baselines. We compare our scheduling framework against the following two baseline methods:
  • Vanilla dense retrieval: the conventional approach that encodes each query and data item as a single vector, corresponding to the “1 Mode” in Eq. 1.
  • POQD [4]: a state-of-the-art multi-vector retrieval framework matching decomposed query with image segments, corresponding to the “1+N Mode” in Eq. 2.
Note that the proposed techniques are designed specifically for retrieval acceleration. Therefore, we do not include the LLM-based query splitting process in our performance comparison.
Datasets. Extending the setup of [4], we evaluate on 4 text-image datasets: CREPE [18], MsCoco [19], NoCaps [20], and Flickr [21], with 1K, 40K, 2K, and 2K images respectively, after data cleaning. Considering the scale of MSCOCO, a subset is randomly sampled for experiments. Note that we treat each image caption as a query and consider a caption relevant only to its corresponding image, following [18].
Models. Embedding and decomposition models are critical to retrieval accuracy. For text-image aligned embedding, we employ CLIP [6] to embed both queries and images. For query decomposition, we use Qwen3-8B [22] as the default large language model, in place of the model used in [4], owing to its improved decomposition performance.
Metrics. To evaluate efficiency, we measure the query throughput (queries per second, QPS) of HiMIR and baselines. For retrieval accuracy, we report the Normalized Discounted Cumulative Gain (NDCG [23]) at top-1 and top-10, denoted as NDCG@1 and NDCG@10.
Figure 6: Trade-off Analysis: Redundant Matching Opt.
VI-B Experiment Results
The main experimental results are summarized in Tab. I. We set four configurations to demonstrate the effectiveness of each proposed techniques: (1) only “1+M+N” algorithm without redundancy reduction. (2) hierarchy pruning (O3), (3) redundant matching optimization(O2), and (4) hierarchy depth optimization (O1). The parameter setting is obtained through the automated configuration framework. Note that we mainly compare our framework with MVR systems (POQD baseline), and thus the average speedup and accuracy improvement of the Vanilla baseline is omitted.
[…]
We further analyze the trade-offs of individual techniques. Fig. 6 illustrates the cross-hierarchy search. The curve shows throughput while the bar shows accuracy. As expected, filtering more images at each granularity (i.e., reducing $T$ and $\alpha$) significantly improves throughput. However, computation does not always translate to accuracy. For example, setting $\alpha$=0.8 yields the highest accuracy on Crepe, surpassing $\alpha$=1.0. It is common in CREPE that some images share common fine-grained objects but differ in composition. Thus hierarchical filtering helps eliminate such interference. This phenomenon is less evident in MS COCO, since Crepe has higher visual complexity. Fig. 7 investigates hierarchy skipping. The trend is consistent across datasets: larger granularity budgets $\tau$ improve accuracy by considering more granularities for finer query–image alignment, at the expense of throughput. This again highlights the intuitive performance–accuracy trade-off.
In summary, although more computation generally favors accuracy, the trade-off between accuracy and performance is non-trivial. This necessitates our automatic configuration framework. As shown in Fig. 8, HiMIR consistently pushes the Pareto frontier. POQD is represented by the green dot at the lower-left corner, while Vanilla dense retrieval lies outside the figure due to its extremely high throughput but poor accuracy.
Figure 7: Trade-off Analysis: Hierarchy Depth Opt.
Figure 8: Pareto Frontier Analysis
VI-C Overhead Analysis
The scheduling overhead of HiMIR consists of two parts. The first is the runtime early-exit metric computation, which compares the ranking of top candidates across consecutive iterations. This step has constant time complexity. The second is the additional sorting per iteration. Since the number of granularities is small (fewer than 10), the overall overhead is negligible, less than 0.1 ms per query.
VII Conclusion
This paper introduced HiMIR, a hierarchical decomposition framework for multi-vector image retrieval. By extending conventional “1+N Mode” retrieval into “1+M+N Mode,” HiMIR adapts to diverse image granularities while systematically exploiting redundancy and sparsity to improve efficiency. Through this co-design of algorithm, computation, and automation, HiMIR achieves significant gains in both accuracy and runtime cost, making fine-grained multimodal retrieval practical for real-world deployment. HiMIR opens up opportunities for broader integration with multimodal LLM systems and for further optimization of specific applications.